OpenStudy (anonymous):

Find the Derivative of f(x)=-6/x at x=12

OpenStudy (anonymous):

@AriPotta Can you help me out on this?

OpenStudy (anonymous):

I can show you guys the work I have so far if you want

OpenStudy (cwrw238):

yes do that

OpenStudy (anonymous):

-6x/x(x+h)-(-6(x+h)/x(x+h))

OpenStudy (anonymous):

I believe I am making a mistake at that step

OpenStudy (cwrw238):

i see - its derivation form first principles

OpenStudy (anonymous):

Yep

OpenStudy (cwrw238):

oops - i'll have to do a little revision on that ...

OpenStudy (anonymous):

From there I would get (-6x-(-6x-6h))/h

OpenStudy (cwrw238):

if no one can help - i;ll be back in about 5-10 minutes

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

I eentually get that it equals 6 but the answer choices are 12, 1/2, 1/24, or 2.

OpenStudy (anonymous):

You guys have anything?

OpenStudy (cwrw238):

ok - i remember now lol! but johnweldon1993 is answering..

OpenStudy (johnweldon1993):

No go right ahead, I'm actually stuck on the last step XD

OpenStudy (johnweldon1993):

Oh I feel stupid XD

OpenStudy (anonymous):

I think I'm gonna go with the wild guess of 24

OpenStudy (johnweldon1993):

Oh the terrible beginning!! lol \[\large \lim_{h -> 0}\frac{f(x + h) - f(x)}{h}\] So since we know f(x) = -6/x...so we plug that in here \[\large \frac{\frac{-6}{x + h} - \frac{-6}{x}}{h}\] We need our common denominator on the top which you found to be x(x + h) \[\large \frac{\frac{-6x - (-6(x + h)}{x(x + h)}}{h}\] \[\large \frac{\frac{\cancel{-6x + 6x} + 6h}{x(x + h)}}{h}\] So we have \[\large \frac{\frac{6h}{x(x + h)}}{h} \] re-write it as \[\large \frac{6h}{hx(x + h)}\] Cancel out your 'h' \[\large \frac{6\cancel{h}}{\cancel{h}x(x + h)}\] \[\large \lim_{h->0}\frac{6}{x^2 + h}\] what does that equal?

OpenStudy (anonymous):

set x=12 and h=0 and you get 1/24

OpenStudy (cwrw238):

good work

OpenStudy (johnweldon1993):

And you are correct!

OpenStudy (anonymous):

I see what I did wrong I took out the common denominator...

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

It wont allow me to close the question...